# Averaging of quantum channels via channel-state duality

### v1

@@ Abstract
Twirling, uniform averaging over symmetry actions, is a standard tool for reducing the description of quantum states and channels to symmetry-invariant data. We develop a framework for averaging quantum channels based on channel-state duality that converts pre- and post-processing averages into a group twirl acting directly on the Choi operator. For arbitrary unitary representations on the input and output spaces, the twirled channel is obtained as an explicit projection onto the commutant of the induced representation on $\mathcal{H}_{\rm out}\otimes\mathcal{H}_{\rm in}$. In the collective setting, where the commutant is the walled Brauer algebra, we introduce a partial-transpose reduction that maps channel twirling to an ordinary Schur-Weyl twirl of the partially transposed Choi operator, enabling formulas in terms of permutation operators. We further extend the construction beyond compact symmetries to reductive non-unitary groups via Cartan decomposition, yielding a weighted sum of invariant-sector projections with weights determined by the Abelian component. Finally, we provide two finite realizations of channel averaging. The first one is a “dual” averaging protocol as a convex mixture of unitary-$1$-design channels on invariant sectors. The second one is a notion of channel $t$-designs induced by weighted group $t$-designs for $t=t_{\rm in}+t_{\rm out}$.
@@ 1 Introduction
Following ideas of [1, 2, 3, 4] we define twirling of quantum channels as follows. Firstly let us introduce notation inspired by previous works on averaging of quantum states [5, 6]. In this work we assume that the dimensions of elementary systems constituting input and output subspaces are equal, however the number of elementary subsystems can be different for input and output spaces. Hence the input and output Hilbert spaces are specifed respectively as: $\mathcal{H}_{\operatorname{in}}=(\mathbb{C}^{d})^{\otimes t_{\operatorname{in}}}$ and $\mathcal{H}_{\operatorname{out}}=(\mathbb{C}^{d})^{\otimes t_{\operatorname{out}}}$. We consider arbitrary channels $\Phi:\mathcal{B}(\mathcal{H}_{\operatorname{in}})\mapsto\mathcal{B}(\mathcal{H}_{\operatorname{out}})$ mapping $t_{\operatorname{in}}$ qu-$d$-its to $t_{\operatorname{out}}$ qu-$d$-its.
@@
Clearly we have: $\mathcal{T}_{U^{-1}}^{\operatorname{in}}(\sigma_{\operatorname{in}})=\pi^{\operatorname{in}}_{\operatorname{U}}(U)^{\dagger}\sigma_{\operatorname{in}}\pi^{\operatorname{in}}_{\operatorname{U}}(U)$, since $\pi^{\operatorname{in}}_{\operatorname{U}}(U^{-1})=\pi^{\operatorname{in}}_{\operatorname{U}}(U)^{-1}=\pi^{\operatorname{in}}_{\operatorname{U}}(U)^{\dagger}$. The last property holds because any representation is a group homomorphism. Finally, let us define twirling of quantum channels as follows (see [1], Sec. II C):
$\mathcal{M}^{\Phi}_{U}(\rho)=\int_{\operatorname{U}(d)}\operatorname{d}\!U\,\mathcal{T}_{U}^{\operatorname{out}}\circ\Phi\circ\mathcal{T}_{U^{-1}}^{\operatorname{in}}(\rho)=\int_{\operatorname{U}(d)}\operatorname{d}\!U\,\mathcal{T}_{U}^{\operatorname{out}}\left(\Phi\left(\mathcal{T}_{U^{-1}}^{\operatorname{in}}(\rho)\right)\right),$ (2)
@@
in which the integration is performed with respect to a Haar measure on the unitary group.
Starting from the above definition of channel twirling as uniform pre- and post-processing by a symmetry representation, we derive an explicit Choi-level expression showing that twirling of channels is equivalent to a group twirl acting directly on the Choi operator under the induced representation $\pi^{\rm out}_{\rm U}\otimes\bar{\pi}^{\rm in}_{\rm U}$, and we express the result as a projection onto the commutant algebra via generalized Schur–Weyl duality. In the collective setting with $t_{\rm in}$ input and $t_{\rm out}$ output subsystems, where the relevant commutant is the walled Brauer algebra, we introduce a partial-transpose reduction that removes the contragredient action and maps the channel twirl to an ordinary Schur-Weyl twirl of the partially transposed Choi operator under $U^{\otimes(t_{\rm in}+t_{\rm out})}$, enabling practical formulas in terms of permutation operators and avoiding an explicit construction of walled Brauer idempotents. Beyond compact symmetries, we extend the approach to averaging over non-unitary operations described by reductive groups using Cartan decomposition and normalized non-compact components, which yields a weighted decomposition of the twirl into invariant-sector projections with coefficients determined solely by the Abelian Cartan part. We also provide two finite realizations that parallel the role of designs in sampling: first, a dual averaging representation in which the (Cartan) channel twirl is implemented as a probabilistic mixture of unitary-$1$-design channels acting on invariant sectors via embedded unitary operator bases, and second, a design-like reconstruction theorem for collective actions showing that any weighted group $t$-design for $t=t_{\rm in}+t_{\rm out}$ induces a channel $t$-design that reproduces the twirled Choi operator from a finite sum.
Main results
Theorem 1 (Channel twirl in Choi form, Section 2).
Let $\Phi:\mathcal{B}(\mathcal{H}_{\rm in})\to\mathcal{B}(\mathcal{H}_{\rm out})$ be a quantum channel with Choi operator $J_{\Phi}=(\Phi\otimes{\rm 1\kern-2.20001ptl}_{\rm{in}})(\rho_{\rm ME})\in\mathcal{B}(\mathcal{H}_{\rm out}\otimes\mathcal{H}_{\rm in})$. Given unitary representations $\pi^{\rm in}_{\rm U}$ on $\mathcal{H}_{\rm in}$ and $\pi^{\rm out}_{\rm U}$ on $\mathcal{H}_{\rm out}$ representing the symmetry transformations on input and output subsystems, define the twirl
$M_{U}^{\Phi}(\rho)=\int_{\mathrm{U}(d)}dU\;T^{\rm out}_{U}\circ\Phi\circ(T^{\rm in}_{U^{-1}})(\rho),\quad T^{\rm in/out}_{U}(X)=\pi_{\rm U}^{\rm in/out}(U)\,X\,\pi_{\rm U}^{\rm in/out}(U)^{\dagger}.$
Then the Choi operator of the twirled channel satisfies
$J_{M_{U}^{\Phi}}=\int_{\mathrm{U}(d)}dU\;\Big(\pi_{\rm U}^{\rm out}(U)\otimes\bar{\pi}_{\rm U}^{\rm in}(U)\Big)\,J_{\Phi}\,\Big(\pi_{\rm U}^{\rm out}(U)\otimes\bar{\pi}_{\rm U}^{\rm in}(U)\Big)^{\dagger},$
i.e. channel twirling is a state twirl on the Choi operator under $\pi_{\rm U}^{\rm out}\otimes\bar{\pi}_{\rm U}^{\rm in}$.
@@
Corollary 1 (Collective unitary twirl reduces to ordinary Schur–Weyl twirl, Section 2).
@@
Assume collective representations $\pi_{\rm U}^{\rm in}(U)=U^{\otimes t_{\rm in}}$ and $\pi_{\rm U}^{\rm out}(U)=U^{\otimes t_{\rm out}}$. Let $\Gamma$ denote partial transpose on the input factor of $\mathcal{H}_{\rm out}\otimes\mathcal{H}_{\rm in}$. Then
@@
$J_{M_{U}^{\Phi}}=\left[\int_{\mathrm{U}(d)}dU\;\Big(U^{\otimes(t_{\rm in}+t_{\rm out})}\Big)\,J_{\Phi}^{\Gamma}\,\Big(U^{\otimes(t_{\rm in}+t_{\rm out})}\Big)^{\dagger}\right]^{\Gamma}.$
@@
In particular, evaluating $M_{U}^{\Phi}$ can be reduced to a standard Schur–Weyl projection of $J_{\Phi}^{\Gamma}$, so one may work directly with permutation operators on $(t_{\rm in}+t_{\rm out})$ tensor factors.
@@
Theorem 2 (Reductive (Cartan) twirl and invariant-sector decomposition, Section 3).
@@
Let $\mathcal{G}$ be a reductive group with Cartan decomposition $G=KAK^{\prime}$, where $G\in\mathcal{G}$, $K\in\mathcal{K}$ and $A\in\mathcal{A}$, $\mathcal{K}$ denotes the maximal compact subgroup of $\mathcal{G}$, $\mathcal{A}$ denotes the maximal Abelian subgroup, and let $\pi_{\mathcal{G}}^{\rm in/out}$ be finite-dimensional matrix representations of $\mathcal{G}$ on $\mathcal{H}_{\rm in/out}$. Define the (normalized) twirl by integrating over $\mathcal{K}\times\mathcal{A}\times\mathcal{K}$ and using the adjoint action on the input side. Then the Choi operator of the twirled channel admits an invariant-sector decomposition
$J_{M_{\mathcal{G}}^{\Phi}}=\sum_{k}\frac{\beta_{k}}{D_{k}}\;\mathcal{P}_{k}(J_{\Phi}),$
@@
where $\{\mathcal{P}_{k}\}$ are the commutant-sector projections determined by the restriction of $\pi_{\mathcal{G}}^{\rm out}\otimes\bar{\pi}_{\mathcal{G}}^{\rm in}$ to the maximal compact subgroup $\mathcal{K}$, while the coefficients $\beta_{k}$ depend only on the Abelian component $\mathcal{A}$ through a real-valued integral. Moreover, in the collective case (e.g. $\mathcal{G}=\mathrm{SL}(d,\mathbb{C})$ with $\mathcal{K}=\mathrm{SU}(d)$ and ${\pi}_{\mathcal{G}}^{\rm in/out}=G^{\otimes t_{\rm in}/t_{\rm out}}$) one may again express $\mathcal{P}_{k}$ using ordinary Schur–Weyl permutation operators after partial transpose.
@@
The Cartan twirl $M_{\mathcal{G}}^{\Phi}$ of the Choi operator can be represented as a probabilistic mixture of unitary-$1$-design channels acting on invariant sectors:
$J_{M_{\mathcal{G}}^{\Phi}}=\sum_{k}\frac{\beta_{k}}{D_{k}}\;\frac{1}{(D_{k}^{G})^{2}}\sum_{l=1}^{(D_{k}^{G})^{2}}\tilde{\gamma}^{(k)}_{l}\,J_{\Phi}\,\tilde{\gamma}^{(k)\dagger}_{l}$,
where $\{\tilde{\gamma}^{(k)}_{l}\}$ are Kraus operators obtained by embedding a unitary operator basis on $\mathbb{C}^{D_{k}^{G}}$ into the full representation space via the duality basis operators.
@@
Let $t=t_{\rm in}+t_{\rm out}$ and suppose $\{(G_{i},w_{i})\}$ is a weighted group $t$-design for a subgroup $\mathcal{G}\subseteq\mathrm{GL}(d,\mathbb{C})$. Then the averaged Choi operator can be reconstructed from a finite sum:
$J_{M_{\mathcal{G}}^{\Phi}}=\left[\int_{\mathcal{G}}dG\;G^{\otimes t}\,J_{\Phi}^{\Gamma}\,(G^{\otimes t})^{\dagger}\right]^{\Gamma}=\left[\sum_{i}w_{i}\;G_{i}^{\otimes t}\,J_{\Phi}^{\Gamma}\,(G_{i}^{\otimes t})^{\dagger}\right]^{\Gamma}.$

### v2

# Choi-level twirling of quantum channels: finite constructions and non-compact transformations

@@ Abstract
Twirling, i.e. averaging over symmetry actions, is a standard tool for reducing quantum states and channels to a symmetry-invariant form. We study channel twirling from the perspective of the channel–state duality and provide a constructive Choi-level description of the averaging map induced by arbitrary input/output representations. Our main technical result concerns the collective setting: for $\pi^{\operatorname{in}}(U)=U^{\otimes t_{\operatorname{in}}}$ and $\pi^{\operatorname{out}}(U)=U^{\otimes t_{\operatorname{out}}}$, we introduce a partial-transpose reduction that removes the contragredient action and converts the mixed (walled Brauer) channel twirl into an ordinary Schur–Weyl twirl of the partially transposed Choi operator under $U^{\otimes(t_{\operatorname{in}}+t_{\operatorname{out}})}$, enabling explicit permutation-based formulas without constructing walled Brauer idempotents or mixed Schur transforms. Beyond compact symmetries, we extend channel twirling to reductive, generally non-unitary groups via Cartan decomposition and obtain an invariant-sector decomposition of the averaged Choi operator with weights determined solely by the Abelian Cartan component. Finally, we provide two finite realizations of channel averaging: a “dual” implementation as a convex mixture of unitary-$1$-design channels acting on invariant sectors, and a design-like reconstruction showing that weighted group $t$-designs induce channel $t$-designs for $t=t_{\operatorname{in}}+t_{\operatorname{out}}$.
@@ 1 Introduction
1.1 General Motivation
Averaging procedures based on symmetry groups play a central role in quantum information theory [1, 2]. For quantum states, uniform averaging (twirling) with respect to a group action provides a systematic way of reducing descriptions to symmetry-invariant data, with applications ranging from entanglement theory to randomized benchmarking and noise-invariant encoding [3, 4, 5, 6, 7, 8]. For quantum channels, averaging is typically defined operationally via uniform pre- and post-processing by symmetry transformations, leading to effective noise models such as depolarizing or stochastic channels [2, 9, 10, 11, 12, 13].
Despite its widespread use, channel averaging is most often treated at the level of operational definitions and specific examples. The Choi–Jamiołkowski isomorphism provides a natural representation of quantum channels as bipartite operators. While this framework is invaluable for tasks such as defining operational distance measures [14] and discriminating quantum operations [15], a general constructive Choi-level formula that works uniformly for arbitrary collective tensor actions and input/output transformations remains largely absent from the literature. Consequently, practical evaluation often relies on case-by-case representation-theoretic machinery. Existing approaches either focus on special classes of symmetries, such as finite or compact groups with irreducible representations, or rely on case-by-case constructions tailored to specific physical settings.
A natural expectation is that channel twirling should correspond to a projection onto symmetry-invariant subspaces of the Choi operator. However, making this correspondence precise for arbitrary input and output transformations, and doing so in a way that is both conceptually transparent and practically usable, turns out to be nontrivial. This difficulty becomes particularly apparent in collective scenarios, where channels act on multiple identical subsystems and the relevant symmetry algebra is no longer given by the symmetric group alone (as in the case of twirling of quantum states) [8], but by the commutant of the tensor product action $U^{\otimes p}\otimes\bar{U}^{\otimes q}$, namely the walled Brauer algebra [16].
1.2 State of the art: covariant and equivariant channels, and mixed Schur–Weyl tools.
Symmetry constraints on quantum channels are most commonly formulated in terms of covariance [^1] It is important to distinguish two related but different notions: covariance/equivariance is a symmetry constraint defining a class of channels, whereas twirling is an averaging procedure that maps an arbitrary channel to a symmetric one.. A channel $\Phi:\mathcal{B}(\mathcal{H}_{\mathrm{in}})\to\mathcal{B}(\mathcal{H}_{\mathrm{out}})$ is called $\mathcal{G}$-covariant if there exist (typically unitary) representations $\pi_{\mathcal{G}}^{\textrm{in}}$ and $\pi_{\mathcal{G}}^{\textrm{out}}$ of the symmetry group $\mathcal{G}$, such that
$\Phi\bigl(\pi_{\mathcal{G}}^{\textrm{in}}(g)\,\rho\,\pi_{\mathcal{G}}^{\textrm{in}}(g)^{\dagger}\bigr)=\pi_{\mathcal{G}}^{\textrm{out}}(g)\,\Phi(\rho)\,\pi_{\mathcal{G}}^{\textrm{out}}(g)^{\dagger}\qquad\forall g\in\mathcal{G}.$ (1)
It is a well-known consequence of the Choi–Jamiołkowski isomorphism that this covariance condition is equivalent to a commutation relation at the Choi level [17, 10, 18]:
$\bigl[\,\pi_{\mathcal{G}}^{\textrm{out}}(g)\otimes\bar{\pi}_{\mathcal{G}}^{\textrm{in}}(g),\,J_{\Phi}\,\bigr]=0\qquad\forall g\in\mathcal{G},$ (2)
that is, the Choi operator must lie in the commutant of the induced representation $\pi_{\mathcal{G}}^{\textrm{out}}\otimes\bar{\pi}_{\mathcal{G}}^{\textrm{in}}$.
When the input and output representations are identical, or related in a direct “same-action” manner (e.g. $U^{\otimes p}$ on the input and $U^{\otimes q}$ on the output for $U\in\mathrm{U}(d)$), the literature often uses the term equivariance to emphasize that applying the symmetry to the input corresponds to applying (essentially) the same symmetry to the output. This distinction is especially natural in the collective setting of many identical subsystems, where one studies channels whose Choi operators commute with the tensor action $U^{\otimes q}\otimes\bar{U}^{\otimes p}$.
A substantial body of work analyzes the structure and applications of such symmetry-restricted channels. For finite groups, a detailed characterization is available in the irreducibly covariant case, where representation-theoretic data yields explicit decompositions of the Choi operator and necessary and sufficient conditions for complete positivity and trace preservation [17, 19]. For continuous unitary symmetries and multipartite settings, mixed Schur–Weyl duality identifies the commutant of $U^{\otimes q}\otimes\bar{U}^{\otimes p}$ with the walled Brauer algebra, providing a powerful language for block diagonalization and parametrization [20, 18, 21]. This perspective underlies modern optimization frameworks: unitary-equivariant semidefinite programs can be reduced to linear programs by expressing feasible operators in a walled Brauer basis and exploiting additional symmetry constraints [22, 18]. On the quantum-algorithmic side, efficient circuit constructions for the mixed Schur transform enable implementations and manipulations of unitary-equivariant channels in a variety of operational settings [20, 23].
In parallel, channel twirling is a standard operational tool, defined by averaging pre- and post-processing operations over a group action [2, 9, 10, 11, 12, 13]. Connections to unitary designs are well understood in several important cases. In particular, twirling by a unitary $1$-design yields a stochastic channel, although the resulting channel may depend on the specific choice of the design [11]. Much of the literature, however, focuses on concrete symmetry models or assumes the invariant decomposition a priori, rather than providing a general and constructive description of channel twirling directly at the level of the Choi operator.
Despite extensive work on covariant channels and mixed Schur–Weyl duality [20, 22, 24, 18, 21, 13], a general and constructive description of channel twirling at the Choi level, especially beyond compact symmetries, has been missing. Moreover in much of the literature on this topic the symmetry is imposed as a structural constraint on $\Phi$ (equivariant/covariant channels), whereas here we focus on the averaging map that sends an arbitrary $\Phi$ to its symmetry-averaged counterpart and we make this map constructive at the Choi level.
1.3 Outline
In this work, we develop a unified framework for averaging quantum channels based on the channel–state duality. Starting from the standard definition of channel twirling as uniform pre- and post-processing by symmetry operations, we show that channel averaging can be expressed exactly as a group twirl acting directly on the Choi operator. More precisely, the Choi operator of a twirled channel is obtained by averaging the original Choi operator with respect to the induced representation on the joint output–input space.
Although this correspondence between operational twirling and a Choi-level group action appears implicitly or explicitly in previous works, we elevate it to a central structural principle. In our approach, the Choi-level formulation is not merely a reformulation, but the fundamental tool that enables a uniform treatment of arbitrary representations of input and output symmetry transformations, collective tensor actions, and non-compact operations.
Viewed in this way, channel twirling becomes an explicit projection onto the commutant of the induced symmetry representation, placing channel averaging exactly within the framework of generalized Schur–Weyl duality [25, 26, 27, 6]. This structural perspective is what allows us to identify invariant-sector decompositions, derive finite-sum formulas, and extend the formalism beyond the compact unitary setting.
In the important collective setting, where the symmetry group acts identically on multiple input and output subsystems, the commutant algebra is given by the walled Brauer algebra. To avoid the technical complexity associated with constructing explicit Brauer idempotents, we introduce a partial-transpose reduction that maps channel twirling to an ordinary Schur–Weyl twirl of a partially transposed Choi operator. As a result, the evaluation of collective channel twirls can be reduced to standard permutation-based formulas, enabling explicit formulas in terms of permutation operators and Schur projectors, without constructing walled Brauer idempotents or mixed Schur transforms.
Genuinely beyond the symmetries represented by compact groups, we extend our framework to averaging over arbitrary finite-dimensional representations of reductive, generally non-compact groups. Using Cartan decomposition, we show that the resulting channel twirl admits a decomposition into invariant-sector projections with weights determined solely by the Abelian component of the group. This provides a controlled and well-defined notion of channel averaging beyond the unitary setting, where naive uniform averaging is not available.
Finally, we address finite realizations of channel averaging. We introduce a dual averaging representation in which the twirled channel is expressed as a convex mixture of unitary $1$-design channels acting on invariant sectors. We further establish a design-like reconstruction result showing that weighted group $t$-designs induce channel $t$-designs for $t=t_{\mathrm{in}}+t_{\mathrm{out}}$, allowing exact reconstruction of the twirled Choi operator from a finite sum.
The results presented here provide a general and constructive description of channel averaging that unifies and extends existing approaches. They offer both conceptual clarification of the role of symmetry in quantum channels and practical tools for analyzing and implementing averaged channels in a wide range of settings.
1.4 Notation
@@
Firstly let us introduce notation inspired by previous works on averaging of quantum states [8, 28]. In this work we assume that the dimensions of elementary systems constituting input and output subspaces are equal, however the number of elementary subsystems can be different for input and output spaces. Hence the input and output Hilbert spaces are specifed respectively as: $\mathcal{H}_{\operatorname{in}}=(\mathbb{C}^{d})^{\otimes t_{\operatorname{in}}}$ and $\mathcal{H}_{\operatorname{out}}=(\mathbb{C}^{d})^{\otimes t_{\operatorname{out}}}$. We consider arbitrary channels $\Phi:\mathcal{B}(\mathcal{H}_{\operatorname{in}})\mapsto\mathcal{B}(\mathcal{H}_{\operatorname{out}})$ mapping $t_{\operatorname{in}}$ qu-$d$-its to $t_{\operatorname{out}}$ qu-$d$-its.
@@
Clearly we have: $\mathcal{T}_{U^{-1}}^{\operatorname{in}}(\sigma_{\operatorname{in}})=\pi^{\operatorname{in}}_{\operatorname{U}}(U)^{\dagger}\sigma_{\operatorname{in}}\pi^{\operatorname{in}}_{\operatorname{U}}(U)$, since $\pi^{\operatorname{in}}_{\operatorname{U}}(U^{-1})=\pi^{\operatorname{in}}_{\operatorname{U}}(U)^{-1}=\pi^{\operatorname{in}}_{\operatorname{U}}(U)^{\dagger}$. The last property holds because any representation is a group homomorphism. Following ideas of [2, 9, 11, 18], let us define twirling of quantum channels as follows:
$\mathcal{M}^{\Phi}_{\operatorname{U}}(\rho)=\int_{\operatorname{U}(d)}\operatorname{d}\!U\,\mathcal{T}_{U}^{\operatorname{out}}\circ\Phi\circ\mathcal{T}_{U^{-1}}^{\operatorname{in}}(\rho)=\int_{\operatorname{U}(d)}\operatorname{d}\!U\,\mathcal{T}_{U}^{\operatorname{out}}\left(\Phi\left(\mathcal{T}_{U^{-1}}^{\operatorname{in}}(\rho)\right)\right),$ (4)
@@
in which the integration is performed with respect to a Haar measure on the unitary group. Further on we will generalize the above averaging procedure to the case of non-compact transformations.
1.5 Main results
Our contributions can be summarized as follows.
  1. 1.
    In the collective setting, we introduce a partial-transpose reduction that removes the contragredient action and converts the mixed Schur–Weyl (walled Brauer) twirl into an ordinary Schur–Weyl twirl on $t_{\operatorname{in}}+t_{\operatorname{out}}$ tensor factors, enabling permutation-based formulas without constructing walled Brauer idempotents.
  2. 2.
    We extend channel twirling beyond compact symmetries to reductive (generally non-compact) groups via Cartan decomposition, obtaining a weighted invariant-sector decomposition with weights determined solely by the Abelian Cartan component.
  3. 3.
    We provide two finite realizations of the channel averaging process: a “dual averaging” implementation as a convex mixture of unitary-$1$-design channels on invariant sectors, and a $t$-design-like reconstruction result showing that weighted group $t$-designs induce channel $t$-designs for $t=t_{\operatorname{in}}+t_{\operatorname{out}}$.
The structural Choi-level identity underpinning channel twirling (Lemma 1) records the standard induced action arising from pre-/post-processing and channel–state duality. Our main technical novelty begins in the collective setting, where Lemma 3 yields a partial-transpose reduction that removes the contragredient action and converts the mixed (walled Brauer) twirl into an ordinary Schur–Weyl twirl on $t_{\operatorname{in}}+t_{\operatorname{out}}$ tensor factors. This enables explicit permutation-operator formulas without constructing walled Brauer idempotents or invoking mixed Schur transforms.
Partial-transpose reduction (collective setting).
@@ 1.5 Main results
Theorem 1 (Collective channel twirl reduces to an ordinary Schur–Weyl twirl, Section 2).
@@
Assume the collective representations $\pi_{\mathrm{U}}^{\operatorname{in}}(U)=U^{\otimes t_{\operatorname{in}}}$ and $\pi_{\mathrm{U}}^{\operatorname{out}}(U)=U^{\otimes t_{\operatorname{out}}}$. Let $\Gamma$ denote partial transpose on the input factor of $\mathcal{H}_{\operatorname{out}}\otimes\mathcal{H}_{\operatorname{in}}$. Then the Choi operator of the Haar-twirled channel satisfies
@@
$J_{\mathcal{M}^{\Phi}_{\mathrm{U}}}=\left[\int_{\operatorname{U}(d)}\mathrm{d}U\;\Big(U^{\otimes(t_{\operatorname{in}}+t_{\operatorname{out}})}\Big)\,J_{\Phi}^{\Gamma}\,\Big(U^{\otimes(t_{\operatorname{in}}+t_{\operatorname{out}})}\Big)^{\dagger}\right]^{\Gamma}.$
@@
Consequently, evaluating $\mathcal{M}^{\Phi}_{\mathrm{U}}$ can be reduced to a standard Schur–Weyl projection of $J_{\Phi}^{\Gamma}$, so one may work directly with permutation operators on $(t_{\operatorname{in}}+t_{\operatorname{out}})$ tensor factors.
Beyond compact symmetries: reductive (Cartan) twirl.
@@
Theorem 2 (Reductive (Cartan) channel twirl and invariant-sector decomposition, Section 3).
@@
Let $\mathcal{G}$ be a reductive group with Cartan decomposition $\mathcal{G}=\mathcal{K}\,\mathcal{A}\,\mathcal{K}$, and let $\pi_{\mathcal{G}}^{\operatorname{in}/\operatorname{out}}$ be finite-dimensional matrix representations on $\mathcal{H}_{\operatorname{in}/\operatorname{out}}$. Define channel twirling by integrating over $\mathcal{K}\times\mathcal{A}\times\mathcal{K}$ (with normalized Abelian component) and using the adjoint action on the input side. Then the averaged Choi operator admits an invariant-sector decomposition
$J_{\mathcal{M}^{\Phi}_{\mathcal{G}}}=\sum_{k}\frac{\beta_{k}}{D_{k}}\;\mathcal{P}_{k}(J_{\Phi}),$
@@
where $\{\mathcal{P}_{k}\}$ are commutant-sector projections determined by the restriction of $\pi_{\mathcal{G}}^{\operatorname{out}}\otimes\bar{\pi}_{\mathcal{G}}^{\operatorname{in}}$ to the maximal compact subgroup $\mathcal{K}$, while the coefficients $\beta_{k}$ depend only on the Abelian component $\mathcal{A}$ via a real-valued integral. Moreover, in the collective case one may again express $\mathcal{P}_{k}$ using ordinary Schur–Weyl permutation operators after applying the partial transpose trick.
Finite realizations: dual averaging and induced channel designs.
@@
The Cartan twirl $\mathcal{M}_{\mathcal{G}}^{\Phi}$ can be represented as a probabilistic mixture of unitary-$1$-design channels acting on invariant sectors:
$J_{\mathcal{M}_{\mathcal{G}}^{\Phi}}=\sum_{k}\frac{\beta_{k}}{D_{k}}\;\frac{1}{(D_{k}^{G})^{2}}\sum_{l=1}^{(D_{k}^{G})^{2}}\tilde{\gamma}^{(k)}_{l}\,J_{\Phi}\,\tilde{\gamma}^{(k)\dagger}_{l}$,
where $\{\tilde{\gamma}^{(k)}_{l}\}$ are Kraus operators obtained by embedding a unitary operator basis on respective irreducible subspaces into the full representation space.
@@
Let $t=t_{\operatorname{in}}+t_{\operatorname{out}}$ and suppose $\{(G_{i},w_{i})\}$ is a weighted group $t$-design for a subgroup $\mathcal{G}\subseteq\mathrm{GL}(d,\mathbb{C})$. Then the averaged Choi operator can be reconstructed from a finite sum:
$J_{\mathcal{M}_{\mathcal{G}}^{\Phi}}=\left[\int_{\mathcal{G}}\mathrm{d}G\;G^{\otimes t}\,J_{\Phi}^{\Gamma}\,(G^{\otimes t})^{\dagger}\right]^{\Gamma}=\left[\sum_{i}w_{i}\;G_{i}^{\otimes t}\,J_{\Phi}^{\Gamma}\,(G_{i}^{\otimes t})^{\dagger}\right]^{\Gamma}.$
Warm-up identity: induced Choi action.
For completeness, we also record the basic Choi-level identity that relates pre/post twirling of channels to the induced action on the Choi operator (Section 2). Since variants of this statement appear in several contexts in the literature, we treat it as a structural lemma underpinning the subsequent constructions.
Lemma 1 (Choi form of channel twirling under an induced representation).
Let $\Phi:\mathcal{B}(\mathcal{H}_{\operatorname{in}})\to\mathcal{B}(\mathcal{H}_{\operatorname{out}})$ be a quantum channel with Choi operator $J_{\Phi}$. Given unitary representations $\pi_{\mathrm{U}}^{\operatorname{in}}$ on $\mathcal{H}_{\operatorname{in}}$ and $\pi_{\mathrm{U}}^{\operatorname{out}}$ on $\mathcal{H}_{\operatorname{out}}$, define
$\mathcal{M}_{\mathrm{U}}^{\Phi}(\rho)=\int_{\operatorname{U}(d)}dU\;\mathcal{T}^{\operatorname{out}}_{U}\circ\Phi\circ\mathcal{T}^{\operatorname{in}}_{U^{-1}}(\rho),\qquad\mathcal{T}^{\operatorname{in}/\operatorname{out}}_{U}(X)=\pi_{\mathrm{U}}^{\operatorname{in}/\operatorname{out}}(U)\,X\,\pi_{\mathrm{U}}^{\operatorname{in}/\operatorname{out}}(U)^{\dagger}.$
Then
$J_{\mathcal{M}_{\mathrm{U}}^{\Phi}}=\int_{\operatorname{U}(d)}dU\;\Big(\pi_{\mathrm{U}}^{\operatorname{out}}(U)\otimes\bar{\pi}_{\mathrm{U}}^{\operatorname{in}}(U)\Big)\,J_{\Phi}\,\Big(\pi_{\mathrm{U}}^{\operatorname{out}}(U)\otimes\bar{\pi}_{\mathrm{U}}^{\operatorname{in}}(U)\Big)^{\dagger}.$